Santorini

Chapters XVI–XVIII of The First Mate’s Log

R. G. Collingwood

Original publication 1940 This edition June 25, 2015 Reformatted May 9, 2017

Edited and annotated by David Pierce

Mathematics Department Mimar Sinan Fine Arts University

http://mat.msgsu.edu.tr/~dpierce/

Contents

Editor’s Preface 3 Bibliography 5

From The First Mate’s Log 6

From Paros to Santorin 7 1 August . . . . . . . . . . . . . . . . . . . . . . . . . 11 2 August . . . . . . . . . . . . . . . . . . . . . . . . . 12

The Monks of the Prophet Elijah 15 3 August . . . . . . . . . . . . . . . . . . . . . . . . . 15 4 August . . . . . . . . . . . . . . . . . . . . . . . . . 21

Monks and Morals 24

Editor’s Preface

Here are chapters xvi, xvii, and xviii of the twenty-two chapters of R. G. Collingwood, The First Mate’s Log : Of a Voyage to Greece in the Schooner Yacht ‘ Fleur de Lys ’ in 1939. The voyagers on the yacht were American, British, and Canadian students from Oxford. Their captain was a twenty-four-year-old American called Chadbourne Gilpatric, who had invited the fifty-year-old Professor Collingwood to join the crew. Collingwood’s account of his life as Gilpatric’s First Mate was published by Oxford in 1940.

The three chapters reproduced here pertain to the yacht’s visit to the island of Thera, also called Santorini. The last of the chapters, ‘ Monks and Morals, ’ also appears as the ninth of the 19 chapters of R. G. Collingwood, Essays in Political Philosophy, edited and with an introduction by David Boucher. Boucher divides the chapters of the collection into two parts. “ Monks and Morals ” is in Part One, which is called “ Political Activity and the Forms of Practical Reason. ” Part Two is “ Civilization and its Enemies. ”

I possess the paperback edition of Boucher’s book, published by Oxford in 1995. It was reading “ Monks and Morals ” there that induced me to look for the whole of The First Mate’s Log. I now have the 2003 reprint published by Thoemmes Press.

For the present document, I extracted the “ Monks and Morals ” chapter from an epub edition of the Essays that I found through Library Genesis on the Web. the editor’s acknowledgment on this chapter reads,

Reprinted from The First Mate’s Log (London and Oxford, OUP and Milford, 1940), 145–53, with the kind permission of Oxford University Press.

I obtained the other two chapters by scanning the pages of the Thoemmes Press edition of The First Mate’s Log and running an ocr program (namely tesseract) on the result. I bracketed the page numbers of that edition with the headings of the pages. In “ Monks and Morals, ” I also give Boucher’s page numbers, in bold. I have reproduced the chart appearing in “ From Paros to Santorin, ” but not the two photographs (which are not very clear in the Thoemmes Press edition anyway). There are no photographs or charts in the other two chapters.

I have read the text for errors, of which I have corrected a few; but I have not checked the text directly against the print version. Possibly the ocr program has missed a comma or confused a semicolon for a colon ; but when I have thought that might be the case, I have been wrong. I have however detected trivial errors in “ Monks and Morals ” as it appears in the Essays; these are indicated in lettered footnotes.

I have attempted to give punctuation marks the spacing that they have in the original. I have done this by replacing those punctuation marks with TEX commands defined appropriately.

Especially in “ Monks and Morals, ” I have added numbered footnotes according to my understanding. Some of these refer to the works of Collingwood listed in the Bibliography immediately following. The first footnote of the “ Monks and Morals ” chapter is Boucher’s. None of the footnotes are by Collingwood. I have also underlined passages that (to me) are noteworthy.

Bibliography

[1] R. G. Collingwood. Speculum Mentis or The Map of Knowledge. Clarendon Press, Oxford, 1924. Reprinted photographically in Great Britain at the University Press, Oxford, 1946.

[2] R. G. Collingwood. The Principles of Art. Oxford University Press, London, Oxford, and New York, paperback edition, 1958. First published 1938.

[3] R. G. Collingwood. Essays in Political Philosophy. Clarendon Press, Oxford, 1989. Edited with an Introduction by David Boucher. Reprinted 2004.

[4] R. G. Collingwood. The Principles of History and other writings in philosophy of history. Oxford, 2001. Edited and with an introduction by W. H. Dray and W. J. van der Dussen.

[5] R. G. Collingwood. The First Mate’s Log: Of a Voyage to Greece on the Schooner Yacht ‘Fleur de Lys’ in 1939. Thoemmes Press, 2003. With an Introduction by Peter Johnson. Reprinted from the 1940 edition.

[6] R. G. Collingwood. An Essay on Philosophical Method. Clarendon Press, Oxford, new edition, 2005. With an Introduction and additional material edited by James Connelly and Giuseppina D’Oro. First edition 1933.

From The First Mate’s Log From Paros to Santorin

Our time of sailing was fixed for two o’clock : but just as we were making sail and heaving short on the anchor officials came on board and desired the Captain to go on shore with them. Half an hour earlier one of the ordinary island steamers (a rather grand one, with two funnels) had come into the harbour ; and we guessed that some busybody on board had thought it his duty to keep the harbour-master up to the mark by instilling into his breast suspicions of ourselves.

This delayed us for an hour. We left under sail, and found the same strong wind blowing from the same quarter outside, with a good lop of sea. We headed into it for the beat to Delos, now dead to windward and fifteen miles away. Not many minutes were needed to make us reconsider our plans. Reefed down as she was, and butting into a short, steep sea, the schooner went to windward slowly and painfully. The wind was too strong to let her carry all plain canvas ; the sea too heavy to let her sail through it without more power than her reefed sails gave her. It was obvious that we should not reach Delos by daylight, and Delos is no place for anchoring in the dark. The Captain called a council of all hands and preferred that we should abandon Delos altogether, and make for Santorin.^{1} The crew agreed unanimously, and we went about.^{2}

1 According to its Wikipedia article (accessed July 24, 2015); the island is classically called Thera, and officially Thira (����); the name Santorini (���������) is a contraction of Santa Irini (St Irene) and was assigned by the Latin Empire of Constantinople (1204–61); the Ottoman name was Santurin or Santoron.

2 The map of Santorin faces page 129.

8  R. G. Collingwood (1889–1943)

[129 : dodging rocks] Immediately to leeward of us were sundry rocks and shoals, just outside the harbour mouth. Beyond these lay the island of Antiparos, with the so-called Fourteen-foot Passage between it and Paros. This passage was our shortest way southward. It is formidably beset by the kind of rocks that you hardly see except in nightmares. Some are like churches with steeples ; some are like half-submerged tortoises ; some are like ruined houses ; some are black, some brown, some white, and some pink ; and against them all the waves were breaking in large spouts of foam. The Fourteenfoot Passage should be easy enough for a navigator furnished with a large-scale chart, a Mediterranean Pilot, and a lead-line. We had them all ; but we had in addition a following wind blowing half a gale, which meant that the lead-line, at least, would be useless. We decided to leave the Fourteen-foot Passage alone and go round the north-west side of Antiparos and Despotikó until we came to Strongylo passage and go through that.

With the wind on the starboard quarter we leaped and hissed between the rocks off the north corner of Antiparos, the Captain standing to windward and addressing them at the top of his voice in terms of opprobrium and contempt whenever they betrayed their whereabouts by spouting foam into the heavens. Some of them we passed within a ship’s length, or so it seemed to the timorous fancy of the First Mate at the wheel ; but we passed them all, and in a few minutes from the conclusion of our council we were out of them and running along the coast of Antiparos, by fantastic cliffs of alternating marble and mica-schist.

From Strongylo passage it is a run of twenty miles [130 : moonlight entry] to the passage between Nio and Sikinos, and at sunset and moonrise, with a sudden drop of the wind, we finished that run and set our course for Santorin sixteen miles away. The wind picked up again and we travelled fast, singing songs. An hour after sunset we sighted the red light of Epáno Meriá, on the northern headland of Santorin ; and at 10.30 we shot through the narrow gateway between Santorin and Therasia, with a strong wind and a rough sea behind us, and all of a sudden found ourselves in the calm waters of the lagoon. The full moon hung ahead of us. Beneath it stood the thousand-foot cliff of the crater-edge, all in deepest shadow. Here and there on the rim of the cliff lights sparkled and the moon shone on something white, the buildings of a village. The largest village, Phirá, four miles away to the southeast, was where we meant to go.

All over the Santorin lagoon, except in a few places on its extreme edge, within a few feet of the cliffs, the water is too deep for anchorage. Even at Phirá, the port of the island, there is no anchorage. There is only a miniature quay to which you can tie up, sidling into it with an anchor dropped close to the cliff to keep your head out. There is also a buoy to which small visiting steamers can be moored. Larger vessels cruise about while they send their boats ashore.

As we approached this quay, a light air fitfully carrying us along, lights began to shine upon it. We must have been seen coming in, and some one had switched the lamps on. Above and behind it loomed the cliffs, with the lights of Phirá twinkling along their tops. Low down in their face white patches, oddly shaped, began to appear. They were rock-cut [131 : inside the crater] dwellings, their fronts neatly whitewashed into the semblance of houses. Close to the quay, and huddled up against the rock-face to northward of it, we could make out a couple of caïques anchored by the bows and warped to the quay by the stern. We sailed up to the nearest of them, sounding all the time ; found bottom only a few yards from the land, swung to port, and dropped our anchor close off her port bow. Heads came anxiously up from her hatchways, but the operation had been neatly timed by the Captain and the wheel skilfully handled by Stephen, and although there was little room for us in the narrow strip of water between the cliff and the place where it became too deep for anchoring, we occupied what room there was without impingeing too closely on our neighbours. It was exactly midnight when we anchored, and after warping her stern to the quay and having a good feed of bread and jam we had turned in by 1.15.

1 August

If the place had been impressive on a midnight entry it was hardly less weird in the light of day. The cliffs were fantastically streaked with lavas and ashes of various colours and fantastically sculptured into ravines and headlands, slopes and bluffs. Round the quay were a few stone houses : farther off on either side the cave-dwellings we had seen overnight. Above, a zigzag stairway led up to the white town.

Men with saddled mules, the local substitute for taxis, waited hopefully opposite the ship. We breakfasted, and then the Captain and First Mate went ashore to see the harbourmaster. We were too proud [132 : expensive water] or too parsimonious (perhaps we already foresaw the future shortage of drachmas) to take mules, and walked up every step of the way, attended by a shady-looking French-speaking parasite and becoming very hot : did our business with the harbourmaster, and had a drink in a café whose balcony overlooked the entire lagoon, with Therasia opposite, and the new lava cones of the Kaïméni islands in between. As usual, the sky was cloudless and the sea deep blue. We could just see the tops of our own masts with the American flag flying at the main. The Captain went back to the ship, leaving me to await the party he should send ashore. I waited for them in that café for three hours, my growing boredom only alleviated by frequent re-reading of the letters I had found at the post office. When they turned up we all went two doors off to have lunch ; and then, in the shop of that excellent photographer Mr. Joakimídes, I spent the afternoon answering my letters.

2 August

All next day, until eight in the evening, I was on harbourwatch duty. There were endless jobs to be done on board, and I had not an idle moment ; not even time to write this log.

We began by taking in water. We put five hundred drachmas’ worth into the sound or port tank. Five hundred drachmas’ worth sounds a lot, but Santorin is a waterless island. There are no springs there, and I do not know of any wells. Winter rain-water is collected on the flat roofs and runs into underground cisterns. When water is needed a board in the floor [133 : short rations] is raised and a small pail on a cord dipped into the cistern beneath.

For purposes of sale the unit is one of those buckets which are used at least from the Mediterranean to the Malay Archipelago, and I do not know how much farther in each direction : the square tin cans of the Standard and other oil companies, with the top out out and a wooden handle inserted. We shipped the fill of 20\frac{1}{2} such tins, estimating the total at 82 gallons, after which our port tank was nearly full. Each tin had to be filled from some cistern or other at the far end of the quay, carried alongside, lowered into a boat, ferried across, lifted on board, and emptied into the tank through a funnel. The First Mate, deeply weighed down by a sense of his responsibility, had to oversee the whole process, keep tally of the tins, and empty each into the funnel with his own hands, wincing in agony whenever a teaspoonful was spilt ; he thought himself a very busy man.

But the water question was not even yet finally cleared up. Santorin was to be our last Greek port, and our next place of call was settled upon as either Malta or some place in Italy, not yet chosen. It would probably take us ten days to reach either port ; and a system of water-rationing had to be imposed. What we ultimately decided was that no fresh water should be used for washing either ship or clothes or persons ; that its use in cooking should be restricted to what should be considered necessary ; and that the ration for drinking should be one pint a day, plus half a pint in the shape of afternoon tea.^{3} This would work out at about half a gallon per man per day, which would be considered a short ration were it not for the fact [134 : the patent log] that we had large quantities of wine on board for drinking at lunch and supper ; more, in fact, than we were at all likely to get through in the time before us.

At first there was a good deal of grumbling over this business of water-rations, combined as it was with a certain degree of rationing in foodstuffs such as jam. One might have fancied that some of us had never in their lives stopped either eating or drinking short of satiety, and believed that no other system offered a life that was worth living. But the grumbling was only half serious, and I will anticipate so far as to say that a week later all hands had settled down to the restrictions with a good enough grace and found them not really intolerable.

3 If this is a British pint, it is 0.568 litre. British or American, two pints make a quart, and four quarts make a gallon.

The Carpenter was another busy man.^{4} The patent log which ticks off the miles of a ship’s progress is ordinarily attached to her by means of a metal bracket on her stern. We had no such bracket and kept our patent log precariously lashed, as I have said, to the gallows. This day the Carpenter made a wooden bracket to which the thing could be lashed ; this bracket itself being attached more or less firmly to one of the fair-leads on the taffrail. But not very firmly ; so a life-line was rigged securing the patent log directly to the gallows in case it should come adrift. It never did.

Minor jobs would be tedious to describe. The day’s chief triumph fell to the Captain and the Second Mate, who between them, with others helping, revolutionized the state of the ship’s rigging. Every wire in it, as I have hinted, had centimetres of slack, and it could not be set up because the rigging screws were rusted fast. This day a concerted attack was made [135 : the rigging-screws] upon them, lasting uninterruptedly for hours together. At last we reaped the fruit of all the lubricating and wrenching that had occupied so large an aggregate of time ever since the cruise began. Every rigging-screw was made to turn and every stay was properly set up.

4 Collingwood was Carpenter as well as First Mate, ‘ and other parts were similarly doubled ’ (page 6 of the original).

The Monks of the Prophet Elijah

3 August

We were all by now much in love with Santorin ; and some of us had found their affections focused upon the monastery of the Prophet Elijah that stands on the top of its highest mountain, and had brought back almost delirious accounts of what they had found there. On this day the Captain and I resolved to see it for ourselves. We set out in the middle of the morning, climbed the stairs on mule-back, and had a light lunch in Phirá.

From Phirá you walk along a loose-surfaced road, made of the dark lava sand which is black when it forms a beach and is wet with sea water, until you reach Pyrgos, five kilometres to the southward. Then from the hill-top of Pyrgos you drop to a low saddle and thence climb the mountain of the Prophet Elijah. We went without haste, for the midday sun is hot in August. In answer to our wishes a small boy suddenly rose out of the ground in front of us, slipped into a vineyard on the left, and came back with two large clusters of grapes, which he solemnly bestowed upon us. Surely the people of Santorin have the prettiest manners in the world. All our crew had similar encounters to narrate : gifts of fruit by the wayside, invitations into strange houses where food and drink were set before them, sweet-smelling herbs offered in [137 : hospitable town] little posies. We thanked our benefactor and ate his grapes as we walked.

Pyrgos was taking its siesta. Hardly a soul was awake. But as we entered the labyrinthine little town we were greeted by an inscription that some one had sprawlingly painted in blue on a whitewashed wall, which said ‘ You are welcome among us ’. After climbing up one side of the town and down the other, through deserted stone-paved streets never more than 4 feet wide, we looked into a courtyard and saw an old lady knitting under a tree with her shoes and stockings beside her chair. It was the only tree we ever saw in Santorin, and we ought, of course, to have expressed surprise and admiration. We were too stupid to do that, but the gambit we did play came off very well. We asked for a drink of water. She called out, and a child came with a bucket.

But this was too important an affair to be left in the hands of a child. From behind the house emerged four strapping young women with aquiline noses like their mother’s and muscular arms. Not beauties ; but the goodwill of their greeting and the eagerness with which they served us proved, if proof was needed, that the message of welcome bestowed by Pyrgos on its visitors was seriously meant. One brought a glass ; one seized the bucket from the little girl and gave it a good rinsing ; and in a minute we were drinking alternate draughts while the girls stood round us in a ring, delighted by our thirst and by their own conviction that the party was a success.

At the monastery we were received by the chaplain, a tall, rosy, well-groomed man of thirty, radiating health and energy ; a man of the world, very much at [138 : the monks] his ease in the presence of strangers. His beard was carefully brushed, his teeth were white, his hands were soft and well kept, and his clothes were exquisite. Chaucer would have revelled in him : ‘ he was a lord ful fat and in good point. ’ There was also a younger monk, a man of very different pattern : ardent, fanatical, downright ; devoid of airs and graces ; tongue-tied in conversation but eloquent, as we were to learn, when his principles were touched. There was a novice of twenty-odd : slender and light of build, fair-haired, his beard still scanty ; rather dog-like in his wistful attentiveness to the menial work that came his way.

There was an old man, massively grey-bearded and raggedly gowned ; his robe had faded to a dull green ; his hands were hard, and he seemed anxious to efface himself, though not less anxious than the rest to see the visitors. There was also a peasant, obviously employed in working the monks’ land, which runs down the mountain on every side in vineyards and wheatfields and tomato-fields from the monastery on the top. It is a great square building, I should fancy of the nineteenth century, and the chapel has stalls for twenty-six monks ; there are now only five.

The fifth was the archimandrite, Papa Loukas, to whose cell we were conducted when those I have mentioned had shown us the chapel. Papa Loukas was a big, grey man of between sixty and seventy. He had the air of authority mixed with modesty, the expression of simplicity combined with thoughtfulness, that become a man who is head of a community, sure of its function in the world and of his ability to rule it wisely in the discharge of that function, and, [139 : the abbot] for himself, wholly devoted to a life that is not a life of study but a life of praise and worship.

Superficially, because the life of praise and worship has a certain likeness to that of study, he reminded me of University scholars I knew. Yet he was not really like them, because he

had a warmth and sweetness of mind which they mostly lack,
and none of that bitter obsession with personal rivalries and
unsolved intellectual problems which mostly disfigures even

the best of them. He had books in his cell, but he did not seem like a man who much read them, nor did they seem like books that were often read. He had grace and dignity of manner, but it was not the manner of one who mixed much with the world, it was the spontaneous manner of a gracious and dignified man. He ordered the novice to make coffee for us, explaining that within the last two days the fast had begun which, lasting until the middle of the month, terminated in the feast of the Panagía, the Assumption of the Blessed Virgin Mary ; and that for this reason the monastery could offer us only the plainest of food : bread, cheese, wine, and tomatoes. This was when we had already asked hospitality for the night, and had been told that we could not be entertained within the monastery, but should sleep in a chapel outside, to which we should go when the gates were locked at sunset.

The bell now began to ring for a four o’clock service, and Papa Loukas led us into the chapel. The service lasted about an hour, and was chiefly devoted to praise of the Panagía. It was sung throughout. The leading part was played by the chaplain, whose dramatic turn of mind gave it an impressive touch of [140 : vespers] pomp. The second monk whom I have called fanatical sang in a lovely tenor, with affecting devotion. The novice took his part with simple piety. But the whole would have been nothing without the archimandrite, whose beautiful bass voice, clear and sweet as a young man’s, filled the painted dome and seemed as though it echoed through caves hidden in the heart of the mountain. The music was elaborate and each man had his own part and knew it without notes ; I suppose it to have been ancient ; certainly it was far superior both in itself and in the fashion of its performance to anything I had previously heard in Greek churches. There was nothing about it either perfunctory or ostentatious. These men were doing their job. They were doing it very well and very conscientiously.^{5} They were not singing their service to impress a congregation ; they were singing it to praise God and the Panagía. Their singing came from the heart, through the mouths of well-trained musicians.

At the end the Captain and I followed the rest round kissing the ikons and crossing ourselves at each one. When we emerged we found an argument in progress. The fanatical monk had protested to the archimandrite not (I suppose) against our admission to the service but against our having been allowed to kiss the ikons. Papa Loukas was dealing with the protest not by a display of authority but by calm and quiet speech. He was explaining, when we came on the scene, that we too were Christians, who had joined in their worship out of piety ; and that it was their duty to welcome us as brothers.

The young monk was overborne ; but not, as it seemed to me, convinced. There was a moment of [141 : evening] awkwardness as we stood in the middle of the discussion about ourselves ; but in a minute it died down and we all went out on the terrace to sit in the sun, to look down at the ruins of the ancient Greek city laid out on its mountain-spur far beneath us, and to hear the names of the villages that we saw disposed up and down the terraced vineyards of the island like places on a map. The sun declined, and all the churches began ringing their bells. Papa Loukas picked out with a loving ear the voice of his own native village Megalokhorió, and beat time to it with one hand as a tear came into his eye. Then our own bells began weaving an ostinato figure in which one bell began and the rest came in one by one until the great bass bell at the top of the tower had made its entry, and the pattern was complete.

5 From the Prologue (page 25) of Collingwood’s Speculum Mentis of 1924 :

there is no truer and more abiding happiness than the knowledge that one

is free to go on doing, day by day, the best work one can do, in the kind

one likes best, and that this work is absorbed by a steady market and thus

supports one’s own life … this freedom and happiness were in principle at

least the lot of every one in the middle ages ; and to what extent they were

actually achieved by no small number of workers we can see when we look

at the work they have left behind them. For these works breathe visibly the

air of a perfect freedom and a perfect happiness. Chaucer and Dante are no

shallow optimists, but their tragedies are discords perpetually resolved in the

harmony of a celestial music.

As the sun went down the shadow of the Prophet’s mountain lengthened itself gigantically, first upon the eastward sea and then against the haze that lay upon it. At last the red disk was gone. We said good-night and withdrew to the little mortuary-chapel a hundred yards beyond the gate. The novice had already carried thither pillows and rugs for our sleep, water and soap and a towel for our hands, and a dinner exactly as Papa Loukas had promised it : chunks of brown bread, small pieces of goat’s-milk cheese, half a dozen tomatoes, and a small bottle of very good wine, the Malmsey that Santorin produces : but also, unpromised, a tiny saucer for each of us with a mouthful of the monastery’s famous rose jam.

When we had eaten it was dusk, and we went to the top of the mountain, which is just north of the monastery wall, to await the moonrise. The west darkened. [142 : moonlight] We sat talking in low voices. There was a haze over the sea ; it was always there, Papa Loukas told us, when the north wind blew ; you saw no horizon ; sea vanished into haze, haze melted into sky. So the moon did not rise, it loomed like a bloodred ghost through the thickness, dark and dim at first, then paling and brightening as it left the haze behind. We turned to go, and there on the top of the wall was Papa Loukas, his head and shoulders outlined against the stars, his cigarette glowing,^{6} and his voice raised in a fatherly benediction.

We slept well on the black-and-white marble floor of our chapel. It must have been about midnight when I awoke and saw Gil standing in the doorway gazing out into the moonlight. The wind had gone. Twenty yards from the door grew a bush covered with pink flowers, a little like hibiscus. I do not know its name. The chapel was full of its sweet smell. Gil turned and lay down again beside me, and we went to sleep.

4 August

At three in the morning we were roused by the bells, hurried into our shoes, and stepped out into the white moonlight and up the rocky path to the monastery. Mass was beginning. It was a longer affair than the service of the afternoon had been ; it lasted two hours, and every minute was beautiful. Taking our cue from the conclusion of yesterday’s argument, we received the consecrated bread, which might have come from the same loaf off which we had dined ; and afterwards, while the sun rose, we were taken to the archimandrite’s cell for coffee and conversation.

[143 : farewells] This took a theologico-political turn. We agreed that no points of doctrine separated the English Church from the Greek ; I am not very strong on the Filioque clause myself, and I was glad to find that Papa Loukas was no more inclined than I was to drag it into the discussion ; and that in matters of Church government and discipline there were no such differences between us that our common detestation of popery could not serve us as an effective bond. We agreed that the Greek and English royal families were intimately connected in a variety of ways, that Greeks and Englishmen were brothers, and that America (which Papa Loukas had visited in his youth) was a great and admirable country. These trite remarks were not really so dull as they seem in the reporting, because in point of fact they were disguised protestations of mutual affection and esteem, the disguise being intended for penetration.

6 So the fast does not include tobacco ?

When talk was of common friends no disguise was needed. Stephen Verney^{7} had spent a night there, and had conversed (intelligibly enough, it appeared) in Balliol Greek. Papa Loukas had taken very kindly to him, and did not mind saying so, until Stephen’s right car must have burned in his sleep. Before six o’clock we had made our farewells, deposited our offering in the chapel chest, and started down the hill, munching the pieces of bread we had saved from our supper.

It was the feast of Metaxas.^{8} All the villages were ringing their bells and working themselves up into a state of obedient jollity. War veterans were wearing their medals. Children male and female were dressed in an imitation of the ungainly costume which in Italy is worn by immature Fascists. No one looked as [144 : gloom] if he or she were really enjoying it. On the way back to Phirá no one offered us grapes. When we got there people were sitting silently about the streets, which had been strewn with leaves, waiting for the fun to begin, and (it seemed) looking forward less eagerly to its beginning than to its end. We hurried down to the ship, and got there just in time for breakfast.

7 Born April 17, 1919, the Right Reverend Stephen Verney had a remarkable obituary in the Guardian on his death on November 9, 2009.

One extract :

In retirement from 1985, Verney was saddened by the pressures for narrowness

and judgmentalism within the Anglican communion, and used his linguistic

skills for a sharp critique of fundamentalism in all faiths.

8 Supported by King George II, General Ioannis Metaxas had seized power in Greece just three years earlier, on August 4, 1936.

Monks and Morals

Most of our crew had visited the monastery,^{9} and none without being charmed and a little awed by what they found there : the atmosphere of earnest and cheerful devotion to a sacred calling, the dignity of the services and beauty of their music, the eager welcome and the loving hospitality, and above all the graces of character and mind which the life either generated in those who had adopted it or at least demanded of aspirants to it and thus focused, as it were, in the place where that life went on.^{10} They were all the more ready to talk about these impressions because they thought them a trifle paradoxical ; they were the least little bit in the world ashamed of such feelings, because to have them seemed a kind of treason to their upbringing.

They had been taught that monks were at worst idle, selfindulgent,^{a} and corrupt ; at best selfishly wrapped up in a wrongheaded endeavour to save their own souls by forsaking the world and cultivating a fugitive and cloistered virtue. They had, I suspect, been taught that the best was worse than the worst ; for whereas a vicious monk was a sinner to be saved, and from another point of view a man doing his best, like most men, to have a good time, a virtuous monk was a man irremediably sunk in the deadliest of moral errors : a man who had renounced the primary duty of helping his fellow men, and had thus corrupted [146 : corruptio optimi ?] the best thing in human nature, the moral principle itself, into the worst, a purely individual and self-centred quest of salvation.^{11}

9 The monastery to which Collingwood refers is that of the Monks of the Prophet Elijah situated near Pýrgos on the Greek island of Santorin (Thira). Collingwood visited the monastery in 1939 during a sea voyage on a schooner yacht crewed by Oxford students and himself. [Note of David Boucher.]

10 See note 5, page 19.

a The word “self-indulgent” is not hyphenated in the epub file; it is hyphenated, but broken between lines at the hyphen, in Boucher’s edition; hyphenated, but unbroken, in the original. Hyphens similarly disappeared when I ran the ocr program on scans of the two earlier chapters.

The music, for example. My friends were not utilitarian in any gross and barely material sense. They would have had nothing but praise for a man who should retire from the world in order to perfect himself in the performance or composition of music. But when that [144] had been done he ought to justify his retirement by coming out of it and making his performance or composition a blessing to his neighbours ; not a means of filling their bellies, but a means of enriching and enlarging their minds. But what could be the social value of music, however beautiful, performed however beautifully to no audience ?

These ideas, if unopposed, would have had no sense of paradox. They would merely have led to disapproval of persons who, as persons, were admittedly charming. The paradox arose because my friends were capable of making judgements not derived from these ideas,^{12} and in this case incompatible with them, and because they had the courage to abide by their judgements. They had seen these monks. They had lived for a time in close converse with them. They had found in them no discernible traces of the moral faults they had been taught to associate with the monkish profession. They had judged them to be good men, although by the standard of social utility they ought to have been bad men ; and on reflection the judgement stood firm.

11 Today the injunction “ Love your neighbor as yourself ” might be seen as having been corrupted into “ Love yourself first. ”

12 Just about everybody would seem to be so capable. They accept what they are taught if it is congenial ; otherwise they reject it. Though he is their spiritual leader, some Catholics will reject Pope Francis’s warn-

Finding myself drawn into discussions of this kind, especially of nights in the cockpit with Dick and Stephen, who now made up the Third Watch, I raised [147 : who is socially useful ?] the question :^{b} How are you to judge social utility ? Granted that social utility is your standard for the value of a man’s life, how is the standard to be applied ? Perhaps it would be reasonable to accept, at least provisionally, the judgement of those neighbours to whom, you say, the man should be useful, as to whether in fact they find him useful. To accept this judgement is to accept the principle that any man is useful to the society in which he lives in so far as his work fulfils a function which, in the opinion of that society, needs for its own welfare to be fulfilled.^{13}

Suppose a man devotes his life to the study of pure mathematics. Is he to be condemned for living on a selfish principle ?

ings about the environmental degradation of the world. Pious Muslims in Turkey have objected to ostentatious displays of wealth by their supposed co-religionists in the dominant political party.

13 The principle is accepted provisionally. Exceptions may appear. In the opinion of addicts, the drug dealer is useful; of students, the teacher who grades easily; of children, the dispenser of free candy. Collingwood will consider exceptions below when he brings up the topic of superstition.

b The period became a colon in Boucher’s edition (both print and electronic).

Not, as my friends readily admitted, on the ground that pure mathematics cannot feed the hungry. Pure mathematics, apart from any consequences which may ultimately come of it, is pursued because it is thought worth pursuing for its own sake. In order to judge its social utility, then, you must judge it not by these consequences but as an end in itself.

What is more, you cannot judge the social utility of a pure mathematician by asking whether he publishes his results.^{14} Unless there is value in being a pure mathematician, there is no value in [145] publishing works on pure mathematics ; for the only positive result these works could have is to make more people into pure mathematicians ; and a society which does not think it a good thing to have one pure mathematician among its members will hardly think it a good thing to have many.

The social justification of pure mathematics as a career in any given society, then, is the fact that the [148 : and who is to judge ?] society in question thinks pure mathematics worth studying : decides that the work of studying pure mathematics is one of the things which it wants to go on, and delegates this function, as somehow necessary for its own intellectual welfare, to a man or group of men who will undertake it. A test for this opinion is that the society in question should be grateful to the pure mathematician for doing his job, and proud of him for being so clever as to be able to do it ; not that every one else should rush in to share his life, but that even if his neighbours feel no call to share it they should honour him for living as he does. The fact that they do so honour him is a proof that they want a life of that kind to be lived among them, and feel its achievements as a benefit to themselves.

14 Rather, you cannot ask only whether he or she publishes. The text becomes a Platonic dialogue here, with Collingwood as Socrates : what he gives us is not necessarily all that he thinks, but what his interlocutors agree to. For myself, inspired in part by other works of Collingwood, I say that mathematics is fundamentally social : it must be communicable to anybody who is interested. It may be that a great mathematician is a poor communicator; but then other mathematicians must step in to do the communicating and so confirm the greatness of the original mathematician. As Collingwood himself says (on page 7) in The Principles of History (which was unfinished at his death in 1943) :

History has this in common with every other science: that the historian is not

allowed to claim any single piece of knowledge, except where he can justify

his claim by exhibiting to himself in the first place, and secondly to anyone

else who is both able and willing to follow his demonstration, the grounds

upon which it is based. This is what was meant, above, by describing history

as inferential.

If this test is applied to the monks of the Prophet Elijah, the answer is not in doubt. Their neighbours are obviously and outspokenly proud of them. You had only to recollect the tone in which people you met asked you, ‘ Are you going to the monastery ? ’ in order to realize that.

But there is an easy way of driving a coach and four through an argument of this kind. You have only to breathe the blessed word ‘ superstition ’. The respectful neighbours are, of course, an ignorant, uneducated, unenlightened, superstitious folk ; the monks are parasites who live on their ignorance and earn through their superstition a respect they do not deserve.

This is a criticism fatal to the line of thought I was suggesting, unless you are prepared to counter-attack it by asking what ‘ superstition ’ is and what part is played in human life by the things to which you give the name. But this would involve a good deal [149 : our enlightened selves ?] of difficult and perhaps humiliating thought. It is simpler to approach the question from another angle, thus :

The assumption underlying this appeal to the idea of ‘ superstition ’ is that the value of a certain man or group of men to

the society in which they live cannot be decided by appeal to

that society’s own judgement, because on this particular question the persons who [146] constitute the society, or whose judgement we refer to when we speak of the society’s judgement, are bad judges, whose verdict is falsified by distorting influences of some kind or other, never mind what, which we call superstitions. But even if these cannot answer the question, the question still arises, and somebody must answer it. Who shall this be ? Plainly the same person who, by using the word ‘ superstition ’, has exploded the claim of his rivals, the persons to whose principles he has applied the word.

The people of Santorin are proud of their monks, and this shows that, according to local opinion, the monks are a valuable section of society, whose work of prayer, and praise, and meditation is work that needs to be done, though not every one has a duty to do it. But, you say, that is only because the people of Santorin are so ignorant and superstitious that they cannot form a true judgement of what needs to be done and what does not. Very good : then it is for us, whose intellectual superiority entitles us to call them ignorant and superstitious, to use that same superiority for doing well what they have done badly. They have judged that their monks are doing a good work, and are of value to themselves. We are not content with their judgement because we think it has been [150 : utility no standard] made on wrong grounds. Let us make up our minds what the right grounds would be, and on those grounds make a judgement of our own.

What would the right grounds be ? Social utility, you say.^{15} Nothing in the activity of one man or class of men is good unless it is useful, for its utility is what constitutes its goodness. I reply that this cannot be true, because it is self-contradictory. An action is useful because it leads to some other action. If this second action is desired only for its utility, that is as much as to say it is desired only because it will lead to a third action. Sooner or later, this series must end ; an action must be reached which is desired not only for its utility but for its own sake : not only because it is expected to lead to something else, but because in itself it is regarded as good.

15 Collingwood will refute social utility as grounds for judging a society.

But it is not clear that he and his interlocutors consider other possibilities.

See below.

If utility is the only goodness, if nothing is good except in so far as it is useful, there is no utility and therefore no goodness : just as, if no commodities had any value except an exchangevalue, none would have even exchange-value, because no exchange would be worth making. To judge all human action by the standard of utility is like establishing a paper currency in which notes can be exchanged only for other notes, never for gold and never for food or drink, tobacco, or railway-tickets, or the services of professional men. To [147] have a currency of that kind is to be bankrupt ; and the same name applies to having only a utility test for the value of human activities.

Sooner or later the judgement that something is good because it is useful rests on the judgement that something is good in itself, irrespectively of whether or not it is also useful. This shilling [151 : moral bankruptcy] is some good to me because it is useful to me, and not for any other reason ; and it is useful to me because by exchanging it I can get things that are good in themselves, things I desire for their own sakes : a fire when I am cold, a meal when I am hungry, or a lift in a bus when I am tired. Nothing is disqualified for being useful by being good in itself ;^{16} a lift, which is good in itself as constituting a rest for weary legs, may also be useful as bringing a man nearer to his destination ; but it is only because some things are good in themselves that anything can be useful.

16 This would seem to hint at the doctrine of the “ overlap of classes ”

The utilitarian trick of judging the worth of all human activities by assessing their utility is therefore logically nonsensical, and hence unworthy of any one who claims to be an educated and enlightened person ; and it is morally disastrous, because it is the first step on the road to a moral bankruptcy brought about by some process in the moral life analogous to inflation in economic life. Inflation pushed to extremity means that real commodities, the things we really want to buy,^{c} cannot be bought ; all we can handle is stuff that is called money ; but nobody wants money, people want the things that money can buy, and if money cannot buy things it forfeits the very name of money. So the moral bankruptcy of which I speak is the experience of finding that life is not worth living, because everything one does is done in the hopes of purchasing by its means a satisfaction which never comes. The way to avoid this moral bankruptcy is to stop judging the value of actions in terms of utility, and to judge them in terms of intrinsic worth.^{17}

developed in An Essay on Philosophical Method of 1933.

c The comma is missing in Boucher’s edition (print and electronic).

Well then, let us try to do it.^{18} Here is Santorin, inhabited by such and such persons whom we have [152 : broadening your mind] met, having such and such manners and customs which we have observed. There is something which we may call the Santorin way of life. It is not a mere aggregate of disconnected units ; it is one pattern into which the monks and the children who gave us grapes and the girls who gave

17 Collingwood speaks of bankruptcy, and life not worth living, in The Principles of Art of 1938 :

I get a certain amount of fun out of writing this book. But I pay for it as I get it, in wretched drudgery when the book goes badly, in seeing the long summer days [95] vanish one by one past my window unused, in knowing that there will be proofs to correct and index to make, and at the end black looks from the people whose toes I am treading on. If I knock off and lie in the garden for a day and read Dorothy Sayers, I get fun out of that too ; but there is nothing to pay. There is only a bill run up, which is handed in next day when I get back to my book with that Monday-morning feeling. Of course, there may be no Monday-morning feeling : I may get back to the book feeling fresh and energetic, with my staleness gone. In that case my day off turned out to be not amusement but recreation. The difference between them consists in the debit or credit effect they produce on the emotional energy available for practical life.

Amusement becomes a danger to practical life when the debt it imposes on these stores of energy is too great to be paid off in the ordinary course of living. When this reaches a point of crisis, practical life, or ‘ real ’ life, becomes emotionally bankrupt ; a state of things which we describe by speaking of its intolerable dullness or calling it a drudgery. A moral disease has set in, whose symptoms are a constant craving for amusement and an inability to take any interest in the affairs of ordinary life, the necessary work of livelihood and social routine. A person in whom the disease has become chronic is a person with a more or less settled conviction that amusement is the only thing that makes life worth living. A society in which the disease is endemic is one in which most people feel some such conviction most of the time.

18 We can speak of intrinsic worth ; but still it is worth only to this or that person. We may admire the life of Santorini after four days ; after four months, we might find it stultifying. Still, if we cannot fully embrace the Santorini lifestyle, there is no reason to think that the islanders would be able to embrace ours. Best would be for everybody to have the

32  R. G. Collingwood (1889–1943)

us water and the unknown person who painted the words of welcome at Pýrgos all fit as parts. What do we think of the Santorin way of life ? Do we think it is a good way or a bad way ? Of course we do not know very much about it ; we have been in the island only four days, and our [148] judgement will very likely be superficial. But we are not asked for a judgement which no further experience could correct. We are asked for a judgement based on the experience we have had : a judgement based on facing the facts that have come our way, facing them with an unprejudiced eye and an open mind.

And if we find ourselves concluding that the Santorin way of life is a good way, and that this commendation of Santorin life as a whole includes the monks who are part of it, we shall know what to do if we find ourselves adding : ‘ All the same as a good protestant I don’t hold with monks ; as a good secularist I don’t think a life of prayer and meditation can be of any value ; and as a good utilitarian I don’t allow myself to praise anything except what I can praise for its utility. ’ We shall reply to ourselves : ‘ What is the use of travel if it doesn’t broaden your mind ? And how can it do that except by showing you the goodness of ways of life which, according to the prejudices you have learned at home, ought to be bad ? ’^{19}

In this way the traveller who began by thinking the men of Santorin ignorant, unenlightened, and super- [153 : socrates once more] stitious may possibly, unless he is very careful, find within his mind a court sitting wherein the men of Santorin rise up in judgement against his own world and against the protestantism, and secularism, and utilitarianism of which it is so proud ; and as judge in that court he may find himself obliged to take their part against his own world ; so that if, later on, his own world should accuse him of not worshipping the -isms that it worships, and of corrupting its young men by imparting to them his heresy, he would have to admit that the accusation was just.

opportunity to embrace another life.

19 In some people, travel seems to confirm the prejudices they left home with! They see what they expect to see.